OpenStudy (anonymous):

Please help not sure if my answer is correct so here is the question... 2x+4 = 13 ----- ---- 3x-3 15 So then I did cross multiply 15x+60=39x-39 -60. -60 15x=39x-99 -39x. -39x --------------- -24x=-99 ---- ---- -24. -24 X=4.125 Is my answer correct? Please help me

OpenStudy (anonymous):

@pooja195

OpenStudy (mathmale):

Suggest that you write 2x+4 = 13 ----- ---- 3x-3 15 as: 2x+4 13 ----- = ---- 3x-3 15

OpenStudy (anonymous):

Yes

OpenStudy (mathmale):

Then 30x+60=? Solve the resulting equation for x. Be certain to check your answer by substituting it into the original equation.

OpenStudy (anonymous):

2 X 15

OpenStudy (anonymous):

@daniel56k how

OpenStudy (mathmale):

Daniel? How could that be? S. Boy is "cross-multiplying."

OpenStudy (anonymous):

Exactly

OpenStudy (mathmale):

Please let S. B. finish his cross multiplication before commenting further. Thanks.

OpenStudy (anonymous):

i mean this 15(2x+4) i was trying to point out his/her mistake

OpenStudy (anonymous):

30x60 @daniel56k

OpenStudy (anonymous):

30x+60

OpenStudy (anonymous):

Then I get X=2!!! :)

OpenStudy (anonymous):

30x+60 is correct 30x+60=39x-39

OpenStudy (anonymous):

Oh we're not done yet lol

OpenStudy (wolf1728):

suicideboy you forgot the second part of the equation

OpenStudy (anonymous):

30x+60=39x-39 now finish this up

OpenStudy (anonymous):

Yay got it

OpenStudy (anonymous):

I got -9x=-21

OpenStudy (wolf1728):

Suicideboymath and x equals ?

OpenStudy (anonymous):

Not sure on my caculator I get

OpenStudy (anonymous):

2.33333

OpenStudy (anonymous):

Or 0.248

OpenStudy (wolf1728):

-9x = -21 is incorrect try solving 30x + 60 = 39x -39 again

OpenStudy (anonymous):

30x-39x=-60-39

OpenStudy (anonymous):

now solve it right

OpenStudy (anonymous):

I'm lost

OpenStudy (wolf1728):

What is 30x -39x ?

OpenStudy (anonymous):

I got -9x/-9= 99/-9

OpenStudy (mathmale):

S. B.: 30x+60=39x-39 Consolidate the x terms on the right side and the constants on the left. First, subtract 30x from both sides. What do you end up with? Next, add 39 to both sides. What does your equation now look like?

OpenStudy (anonymous):

So X=-11!!!! :D

OpenStudy (anonymous):

-60-39=-99 not 99

OpenStudy (mathmale):

Check x=-11! Check it! How would you do that?

OpenStudy (anonymous):

+11

OpenStudy (anonymous):

X=11

OpenStudy (anonymous):

now correct the mistake here -9x/-9= 99/-9

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

11 is correct

OpenStudy (anonymous):

Yay thanks Danny :)

OpenStudy (anonymous):

you are welcome